OpenStudy (anonymous):

1. Which equation represents a nonlinear function? A) y = x + 7 B) y = 3x - 1 C) y = 3x^2 + 1 D) y = x - 1/2

OpenStudy (accessdenied):

Have you seen this formula for slope-intercept of a line? y = mx + b

OpenStudy (accessdenied):

Of the four options, three are in this format of a line. The leftover has a foreign exponent on the x that nullifies it from linear-dom.

OpenStudy (anonymous):

Yes I have seen the formula for the slope.

OpenStudy (accessdenied):

The key is that A, B, and D are all in the form of a linear equation y = mx + b. C, however, seems to have that x^2. That doesn't appear in the slope-intercept form, right?

OpenStudy (anonymous):

Yes.

OpenStudy (accessdenied):

So, it does not meet the qualifications of a linear equation. That would be what you are looking for! :)

OpenStudy (anonymous):

Lol okay.

OpenStudy (accessdenied):

In general, a linear equation only has x and y to the first power and no funny business with x and y multiplied together. No special functions, either. Glad to help!

OpenStudy (anonymous):

Thanks alot :)

OpenStudy (anonymous):

2. Which is NOT true about the above table? A) It shows a linear function. B) The variable x increases by 1 each time. C) The rate of change is not constant. D) The variable y increases by a different value each time.

OpenStudy (kohai):

Do you notice a pattern about what happens to the y value in relation to the x?

OpenStudy (nincompoop):

y = x^2 ?

OpenStudy (kohai):

Very good Nin :D

OpenStudy (anonymous):

So it is D?

OpenStudy (kohai):

I believe it would be A. This graph would create a parabola, not a linear one. In terms of D, y is changing by a different value each time.

OpenStudy (anonymous):

So A not D right because it doesn't show a linear function.

OpenStudy (kohai):

That's correct. A linear function would make a linear graph.

OpenStudy (anonymous):

Lol thanks so much guys for helping me understand this!

OpenStudy (kohai):

Good luck :)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Last question is Wrong.

OpenStudy (anonymous):

???? then what is the correct answer

OpenStudy (anonymous):

it would be c, right ? the rate of change is not constant